A solenoid has a magnetic reluctance of 0.0022. It has 300 turns. If the current flowing is 0.5 A, find the total flux in max wells.


1
Expert's answer
2022-03-28T14:25:21-0400

Reluctance

"r=\frac{I}{\mu_0 A}"

"A=\frac{I}{\mu_0 r}"


"A=\frac{0.5}{4\times3.14\times10^{-7}\times0.0022}=18.09\times\times10^{7}m^2"

Magnetic field

"B=\mu_0 ni"


"B=4\times3.14\times10^{-7}\times300\times0.5=1.884\times10^{-4}T"

Magnetic flux


"\phi=BA=1.884\times10^{-4}\times18.09\times10^{7}=34.081\times10^{3}T"